A sample of magnesium carbonate is found to contain 1.39x10^24 atoms of oxygen. How much does the sample weigh?
1
Expert's answer
2017-11-28T15:28:06-0500
In 1 mole of magnesium carbonate (MgCO3) there’re 3 moles of Oxygen.
Also it is known that 1 mole of substance contains 6.02×1023 particles.
Find the amount of oxygen in the sample:
ν = 〖1.39×10〗^24/〖6.02×10〗^23 = 2.3 (mole)
Find the amount of magnesium carbonate in the sample:
ν = 2.3/3 = 0.77 (mole)
Find the mass of the sample (M(MgCO3) = 84 (g/mole)):
m = 84×0.77 = 64.68 (g)
Answer
The sample weighs 64.68 g.